OpenStudy (anonymous):

An average of 50000 people visit zoo per day. The zoo charges $12 for admission. It is predicted that for each $1 increase the zoo would lose 1500 customers a day. Express daily revenue from ticket sales,R, as a function of the number of $1 price increases,x

OpenStudy (anonymous):

Total money=50000*12 total money (with change)= (50000*12)-(1500*13)

OpenStudy (anonymous):

im mistaken, that second equation should be: total=(50000*13)-(1500*13)

OpenStudy (anonymous):

x should represent the 1 in both the increase in revenue and price and decrease in customers

OpenStudy (anonymous):

Thank you guys...

OpenStudy (anonymous):

I'm not liking that answer though it doesn't graph right

OpenStudy (anonymous):

R(x)=(50000(12+x))-1500x(12+x) @Oudi90 This graphs properly, missed the x after the 1500 the first time.

OpenStudy (anonymous):

maximum revenue will be realized at x=10.66 ($22.66 per ticket) revenue at $770,666.67 and attendance at 34,010 daily visitors.

OpenStudy (anonymous):

Much better thank ou ehuman